# CoordField: Coordination Field for Agentic UAV Task Allocation In Low-altitude Urban Scenarios

### v3

@@ Abstract
With the increasing demand for heterogeneous Unmanned Aerial Vehicle (UAV) swarms to perform complex tasks in urban environments, system design now faces major challenges, including efficient semantic understanding, flexible task planning, and the ability to dynamically adjust coordination strategies in response to evolving environmental conditions and continuously changing task requirements. To address the limitations of existing approaches, this paper proposes coordination field agentic system for coordinating heterogeneous UAV swarms in complex urban scenarios. In this system, large language models (LLMs) is responsible for interpreting high-level human instructions and converting them into executable commands for the UAV swarms, such as patrol and target tracking. Subsequently, a Coordination field mechanism is proposed to guide UAV motion and task selection, enabling decentralized and adaptive allocation of emergent tasks. A total of 50 rounds of comparative testing were conducted across different models in a 2D simulation space to evaluate their performance. Experimental results demonstrate that the proposed system achieves superior performance in terms of task coverage, response time, and adaptability to dynamic changes.
@@ I Introduction
The growing demand for deploying Unmanned Aerial Vehicle (UAV) swarms to perform tasks such as pedestrian detection, vehicle tracking, and traffic signal monitoring in complex and dynamic urban environments has highlighted the advantages of multi-UAV systems over single UAV, particularly in terms of coverage, system redundancy, and operational efficiency. However, the increasing complexity of tasks and the highly dynamic nature of urban environments also pose significant challenges. Effective task allocation in such settings requires the systems to continuously reassign UAVs to new tasks, adapt to UAV failures and changing task priorities, and achieve decentralized intelligent coordination without relying on centralized control.
To address these challenges, various optimization algorithms have been developed for UAV swarm coordination, such as the Grey Wolf Optimizer [1] and the Whale Optimization Algorithm [2]. However, these methods are typically tailored to specific types of tasks and struggle to adapt to heterogeneous UAV clusters operating in highly dynamic environments. With the rise of large language models (LLMs), recent research has explored leveraging their powerful semantic understanding and tool-calling capabilities to address such problems. For example, several approaches attempt to utilize few-shot or zero-shot learning methods based on existing models, enabling LLMs to perform multi-round planning, itinerary calibration, and execution [3, 4]. Nevertheless, such methods still exhibit significant limitations when handling multiple tasks in complex and rapidly changing environments.
@@ II-A Agentic Systems
In recent years, Agentic AI has emerged as a novel intelligent paradigm characterized by autonomous perception, goal-driven planning, and adaptive decision-making, offering innovative and effective solutions for multi-agent collaborative tasks [8, 9]. Unlike traditional passive approaches, Agentic AI emphasizes continuous goal pursuit, contextual memory, adaptive tool use, and self-reflection capabilities of agents [10]. These abilities enable agents to operate with greater autonomy and coordination in dynamic environments. Agent systems powered by LLMs have demonstrated strong reasoning and planning capabilities, particularly in complex task decomposition and sequential execution [3]. These features provide a solid theoretical and methodological foundation for building autonomous multi-UAV systems. In complex urban task scenarios [11], Agentic AI can assist UAVs in understanding high-level instructions, distributing subtasks effectively, and rapidly reconfiguring coordination strategies in response to task changes or agent failures, thereby enabling truly autonomous collaboration [12, 13].
@@ II-B LLMs for Natural Language Task Parsing
With the rapid advancement of LLMs in natural language understanding and general reasoning, LLMs have demonstrated superior capabilities over traditional rule-based parsers in handling ambiguous semantics, conditional logic, and spatial language. They also exhibit strong generalization in task generation, even under limited domain-specific data. Researchers have developed several frameworks such as Code-as-Policies and Prompt2Action[14, 15], enabling LLMs to map natural language instructions into task-executable codes. These approaches often incorporate few-shot prompting and constraint prompting to improve the validity and safety of the generated outputs. Furthermore, the GSCE framework, which consists of Guidelines, Skill APIs, Constraints, and Examples, has demonstrated the practical value of prompt engineering in UAV task planning[16], guiding LLMs to generate physically feasible and constraint-compliant control programs by providing skill APIs, operational constraints, and illustrative examples.
@@ II-C Field-Based Coordination and Allocation
Field-driven methods, particularly the Artificial Potential Field (APF) approach, have been widely applied in the field of mobile robotics[17]. In this approach, target points are modeled as attractive sources, obstacles as repulsive sources, and robots navigate by descending along the gradient of the resulting potential field, thereby enabling path planning and real-time obstacle avoidance. This approach has been applied applications in traffic flow modeling [18], where field-based methods are used to simulate the evolution of traffic dynamics. Vehicles are modeled as particles influenced by traffic density fields, velocity fields, and other factors, allowing the system to capture both macroscopic and microscopic variations more effectively [19].
Inspired by this line of thought, our work adopts a similar strategy by modeling task demand and UAV workload as continuous spatial fields. A dynamically evolving potential field is constructed based on the distribution of task urgency, guiding UAVs to move and allocate themselves through local gradient descent [20]. The advantage of this method lies in its decentralized structure, strong real-time adaptability, and high responsiveness to task changes, making it particularly well-suited for urban scenarios where tasks are unevenly distributed and dynamically changing. Compared with traditional discrete optimization approaches, field-driven methods offer a more flexible and scalable solution.
@@ IV-A Environment and UAV Configuration
In this environment, UAVs operate within the airspace above the road network and maintain a consistent, safe flight altitude to simplify operations and ensure task execution uniformity. A total of 20 UAVs are deployed in the simulation, divided into two types: patrol UAVs and tracking UAVs, with 10 units each. While both types share identical flight performance characteristics, they differ in sensing and recognition capabilities. Patrol UAVs are better suited for wide-area scanning and anomaly detection tasks, whereas tracking UAVs are designed for precise following of specific targets. Each UAV is assigned a capability vector $\mathbf{c}_{i}(t)$, which dynamically represents its adaptability to different types of tasks.
To enhance the dynamism and complexity of the simulation, the environment is updated in real time with a time step $T$, allowing accurate modeling of pedestrian and vehicle movements, as well as changes in traffic signal states. Task instructions are issued via natural language input, such as “Please inspect the crowd and vehicles.” These instructions are parsed through the DeepSeek API, which interprets the natural language input and transforms it into structured task information, including parameters such as target location, task type, and priority.
@@ IV-B Experimental Results and Analysis
Fig. 4 presents the optimized UAV trajectories. Each blue trajectory connects a starting point (green triangle) to a task target (red circle), demonstrating the effect of global path optimization based on minimum travel distance. The results show that the UAVs efficiently cover multiple task points while maintaining a reasonable spatial distribution, thereby avoiding congestion and conflict.
The experimental evaluation metrics include the following 4 aspects:
1) Task Parsing Accuracy (TPA): This metric measures the proportion of natural language instructions that are correctly parsed into structured task representations. It reflects the system’s ability to interpret high-level human commands, which is critical for enabling intuitive human–agent interaction and ensuring accurate downstream execution.
2) Coverage Efficiency (CE): This quantifies the degree to which UAVs effectively cover the areas defined by the task potential field $\phi(x,y,t)$. It is computed as the ratio between the UAV-covered region and the total task-defined region. High coverage efficiency indicates the system’s effectiveness in spatial distribution and responsiveness to task urgency.
3) Task Load Balance (TLB): This evaluates how evenly the task load is distributed across all UAVs, using the norm of the task allocation density gradient $\nabla\rho$. A balanced load prevents overutilization of individual UAVs, reduces task redundancy, and ensures system robustness in dynamic environments.
4) UAV Utilization Rate (UUR): This metric represents the proportion of UAVs actively engaged in task execution relative to the total available fleet. Higher utilization reflects efficient agent coordination and resource deployment, which is essential for maximizing mission effectiveness and operational economy.
@@
The experimental evaluation on 50 natural language command tasks demonstrates that the proposed coordination field agentic system outperforms all baseline models across multiple key performance metrics as shown in Table 1. Specifically, it achieves the highest TPA of 96%, significantly exceeding GPT-4o (74%), Claude-3-7-Sonnet (76%), and Deepseek-v3 (70%), indicating its superior semantic understanding and robustness in converting unstructured commands into executable UAV instructions. In terms of CE, the system maintains an average rate of 95%, consistently outperforming other models such as GPT-4o (85%) and LLaMA-4-Scout (79%), and remains highly reliable even under task-burst scenarios with over 80% effective coverage. With respect to TLB, our method demonstrates the lowest task distribution deviation of only 0.8, compared to 2.1 and 1.6 observed in Deepseek-v3 and Claude-3-7-Sonnet respectively, which significantly reduces resource clustering and enhances spatial coordination. Additionally, the UUR reaches 97%, confirming that the majority of UAVs remain actively engaged in task execution, whereas competing models such as GPT-4o and LLaMA-4-Scout fall below 85%. These results collectively validate the system’s strong coordination capability, dynamic adaptability, and overall efficiency in managing heterogeneous UAV swarms within complex urban environments.
@@
This paper proposes a coordination field Agentic system for UAV task coordination. The system demonstrates high efficiency, stability and adaptability, and is able to parse natural language instructions in real time and achieve efficient task allocation through coordination field guidance. Experimental results verify its superior performance in complex and dynamic urban environments. In future work, we plan to extend the system to three-dimensional scenarios and real-world UAV swarms[22, 23].

### v4

@@ Abstract
With the increasing demand for heterogeneous Unmanned Aerial Vehicle (UAV) swarms to perform complex tasks in urban environments, system design now faces major challenges, including efficient semantic understanding, flexible task planning, and the ability to dynamically adjust coordination strategies in response to evolving environmental conditions and continuously changing task requirements. To address the limitations of existing methods, this paper proposes CoordField, a coordination field agent system for coordinating heterogeneous drone swarms in complex urban scenarios. In this system, large language models (LLMs) is responsible for interpreting high-level human instructions and converting them into executable commands for the UAV swarms, such as patrol and target tracking. Subsequently, a Coordination field mechanism is proposed to guide UAV motion and task selection, enabling decentralized and adaptive allocation of emergent tasks. A total of 50 rounds of comparative testing were conducted across different models in a 2D simulation space to evaluate their performance. Experimental results demonstrate that the proposed system achieves superior performance in terms of task coverage, response time, and adaptability to dynamic changes.
@@ I Introduction
The growing demand for deploying Unmanned Aerial Vehicle (UAV) swarms to perform tasks such as pedestrian detection, vehicle tracking, and traffic signal monitoring in complex and dynamic urban environments has highlighted the advantages of multi-UAV systems over single UAV, particularly in terms of coverage, system redundancy, and operational efficiency. However, the increasing complexity of tasks and the highly dynamic nature of urban environments also pose significant challenges[1]. This requires effective task allocation methods for drone swarms to adapt to ever-changing new tasks, while also achieving real-time intelligent scheduling of the system.
To address these challenges, various optimization algorithms have been developed for UAV swarm coordination, such as the Grey Wolf Optimizer [2] and the Whale Optimization Algorithm [3]. However, these methods are typically tailored to specific types of tasks and struggle to adapt to heterogeneous UAV clusters operating in highly dynamic environments [4]. With the rise of large language models (LLMs), recent research has explored leveraging their powerful semantic understanding and tool-calling capabilities to address such problems. For example, several approaches attempt to utilize few-shot or zero-shot learning methods based on existing models, enabling LLMs to perform multi-round planning, itinerary calibration, and execution [5, 6]. Nevertheless, such methods still exhibit significant limitations when handling multiple tasks in complex and rapidly changing environments.
@@ II-A Agentic Systems
In recent years, Agentic AI has emerged as a novel intelligent paradigm characterized by autonomous perception, goal-driven planning, and adaptive decision-making, offering innovative and effective solutions for multi-agent collaborative tasks [10, 11]. Unlike traditional passive approaches, Agentic AI emphasizes continuous goal pursuit, contextual memory, adaptive tool use, and self-reflection capabilities of agents. These abilities enable agents to operate with greater autonomy and coordination in dynamic environments. Agent systems powered by LLMs have demonstrated strong reasoning and planning capabilities, particularly in complex task decomposition and sequential execution [5]. These features provide a solid theoretical and methodological foundation for building autonomous multi-UAV systems. In complex urban task scenarios [12], Agentic AI can assist UAVs in understanding high-level instructions, distributing subtasks effectively, and rapidly reconfiguring coordination strategies in response to task changes or agent failures, thereby enabling truly autonomous collaboration [13, 14].
@@ II-B LLMs for Natural Language Task Parsing
With the rapid advancement of LLMs in natural language understanding and general reasoning, LLMs have demonstrated superior capabilities over traditional rule-based parsers in handling ambiguous semantics, conditional logic, and spatial language[15]. They also exhibit strong generalization in task generation, even under limited domain-specific data. Researchers have developed several frameworks such as Code-as-Policies and Prompt2Action[16, 17], enabling LLMs to map natural language instructions into task-executable codes. These approaches often incorporate few-shot prompting and constraint prompting to improve the validity and safety of the generated outputs. Furthermore, the GSCE framework, which consists of Guidelines, Skill APIs, Constraints, and Examples, has demonstrated the practical value of prompt engineering in UAV task planning[18], guiding LLMs to generate physically feasible and constraint-compliant control programs by providing skill APIs, operational constraints, and illustrative examples[19, 20].
@@ II-C Field-Based Coordination and Allocation
Field-driven methods, particularly the Artificial Potential Field (APF) approach, have been widely applied in the field of mobile robotics[21]. In this approach, robots navigate by descending along the gradient of the resulting potential field, thereby enabling path planning and real-time obstacle avoidance. This approach has been applied applications in traffic flow modeling [22], where field-based methods are used to simulate the evolution of traffic dynamics. Vehicles are modeled as particles influenced by traffic density fields, velocity fields, and other factors, allowing the system to capture both macroscopic and microscopic variations more effectively .
Inspired by this line of thought, our work adopts a similar strategy by modeling task demand and UAV workload as continuous spatial fields. A dynamically evolving potential field is constructed based on the distribution of task urgency, guiding UAVs to move and allocate themselves through local gradient descent [23]. This method excels with its decentralized structure, real-time adaptability, and responsiveness to dynamic tasks, making it ideal for uneven and changing urban scenarios. Unlike traditional discrete optimization, field-driven approaches are more flexible and scalable.
@@ IV-A Environment and UAV Configuration
In the simulation, 20 UAVs operate at a fixed, safe altitude above a $1000\times 1000$ urban road network. They are evenly divided into patrol and tracking types—sharing identical flight dynamics but differing in sensing capabilities, with patrol UAVs suited for wide-area scanning and tracking UAVs for precise target following. Each UAV is assigned a dynamic capability vector $\mathbf{c}_{i}(t)$ to reflect task adaptability. The environment updates in real time with step $T$, capturing pedestrian and vehicle dynamics and traffic signal changes. Task commands, such as “Please inspect the crowd and vehicles,” are issued via natural language and parsed by the DeepSeek API into structured task parameters including location, type, and priority.
@@ IV-B Experimental Results and Analysis
Fig. 4 presents the optimized UAV trajectories. Each blue track represents a mission route, with a total of ten, starting from the starting point (green dot) to the task target (red star), showing the global path optimization effect based on the minimum travel distance while maintaining a reasonable spatial distribution to avoid congestion and conflict.
In order to fully verify the performance of the proposed CoordField system, this study designed comparative experiments from two perspectives: on the one hand, comparing the differences in task understanding and execution processes of different LLMs; on the other hand, while ensuring that other parts of the agentic system remain unchanged, further investigating the differences in task performance between different traditional task allocation algorithms and the Coordination field. The following is a detailed analysis.
@@
    TPA: Task Parsing Accuracy; CE: Coverage Efficiency; TLB: Task Load Balance; UUR: UAV Utilization Rate.
The experiment in Table I compares the performance of the CoordField system with the current mainstream large language models, in natural language instruction parsing and task collaborative execution. The experimental results show that CoordField has achieved leading performance in all indicators from four dimensions. The multi-agent system designed by the CoordField system can gain advantages in language understanding, multi-task parsing, and resource scheduling, especially in multi-round instruction scenarios, showing stronger generalization and robustness.
TABLE II: Comparison of performance with traditional algorithms
Algorithm CR CE TLB UUR
CoordField 89% 95% 0.8 97%
ACO 93% 84% 1.3 68%
GWO 91% 89% 1.7 75%
WOA 97% 90% 1.7 91%
A* 93% 75% 3.6 53%
  • ACO: Ant Colony Optimization; GWO: Grey Wolf Optimizer; WOA: Whale Optimization Algorithm; A: A Star Algorithm.
    CR: Completion Rate; CE: Coverage Efficiency; TLB: Task Load Balance; UUR: UAV Utilization Rate.
The comparative experiment in Table II evaluates the performance of the Coordination field method by replacing the Coordination field part of the system with four other classic optimization algorithms while ensuring that the entire Agentic system structure remains unchanged. From the experimental data, it can be seen that although the completion rate is slightly better than CoordField, reflecting the accuracy of the traditional optimization algorithm, CoordField performs best in terms of task coverage efficiency and load balance. In contrast, the traditional optimization method has large fluctuations in balance. This group of experiments verifies that the Coordination field method introduced by CoordField can achieve a more efficient multi-UAV collaborative allocation strategy.
@@
This paper proposes a coordination field Agentic system for UAV task coordination. The system demonstrates high efficiency, stability and adaptability, and is able to parse natural language instructions in real time and achieve efficient task allocation through coordination field guidance. Experimental results verify its superior performance in complex and dynamic urban environments. In future work, we plan to extend the system to three-dimensional scenarios and real-world UAV swarms.